OpenStudy (anonymous):

HELP PLEASE? find the indefinite integral. ∫ cos 6x dx

OpenStudy (anonymous):

Most people do that by substitution, however I'd recommend you to get used to several guess methods considering such integrals. Because I am pretty sure you know how to integrate: \[\Large \int\cos(x)dx \]

OpenStudy (anonymous):

@Spacelimbus okay

OpenStudy (anonymous):

So you can guess that :\[\Large \int \cos(6x)dx\approx \sin(6x)+C \] But that's not quite correct yet is it? Apply the chain rule and you'll see that you get an unwanted 6 infront of the cos(6x) term. So, you only need to fix that.

OpenStudy (anonymous):

okay thanks (:

OpenStudy (anonymous):

@Spacelimbus so it would be

OpenStudy (anonymous):